# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ 2.3 A graph-minor analogue of Dyck’s lemma
Let $k,h,c\geq 0$ be integers such that $k\geq 1$ and $c\geq 3.$ Let $H^{\prime}$ be the $(18k,4(c+h+1)\cdot 18k)$-cylindrical grid and $H^{\prime\prime}$ be the $(18k,4(c-2+h+1+1)\cdot 18k)$-cylindrical grid. Let $I^{\prime}_{c},I^{\prime}_{h}$ be a partition of $[2,c+h+1]$ where $|I^{\prime}_{c}|=c$ and $|I^{\prime}_{h}|=h$ and there exists some integer $i\in I^{\prime}_{c}$ such that $i+1,i+2\in I^{\prime}_{c},$ moreover, let $I^{\prime\prime}_{c},I^{\prime\prime}_{h}$ be a partition of $[2,c+h]$ such that $i\in I^{\prime\prime}_{h}$ and $i+1\in I^{\prime\prime}_{c}.$ Finally, let $H^{1}$ be the $(k,4(c+h)\cdot k)$-cylindrical grid and $H^{2}$ be the $(k,4(c+h+1)\cdot k)$-cylindrical grid.
@@
    the mixed surface grid $H^{\prime}_{I^{\prime}_{c},I^{\prime}_{h}}$ of order $18k$ with $c$ crosscaps and $h$ handles contains the mixed surface grid $H^{1}_{I^{\prime\prime}_{c},I^{\prime\prime}_{h}}$ of order $k$ with $c-2$ crosscaps and $h+1$ handles as a minor, and
@@
    the mixed surface grid $H^{\prime\prime}_{I^{\prime\prime}_{c},I^{\prime\prime}_{h}}$ of order $18k$ with $c-2$ crosscaps and $h+1$ handles contains the mixed surface grid $H^{2}_{I^{\prime}_{c},I^{\prime}_{h}}$ of order $k$ with $c$ crosscaps and $h$ handles as a minor.
@@
The final step is to untangle the crosscap at position $i+2$ from the newly created handle. Notice that “position $i+2$” refers only to the bundle of paths forming the crosscap transactions as we are not in a mixed surface grid any more. Indeed, by merging the other two crosscaps into a single handle, we have lost a position. Similar to the proof of Subsection 2.2 we fix two disjoint intervals of endpoints, one within the right most part of the crosscap transaction on the cycle $C_{18k},$ and the other in the right most part of the middle transaction that is part of the handle. Both intervals are of order $2k.$ We then route a crosscap transaction from the left of these intervals to the right. As depicted in Figure 6 we may traverse the handle transaction which allows us to traverse the crosscap exactly once. Hence, the resulting transaction is, indeed, a crosscap. Since each of the two transactions involved in the handle are of order $12k,$ and we only spend $4k$ of their paths to route the crosscap, there are enough paths left to choose $k$ paths each in order to form a handle transaction of order $k.$ By restricting all other handles and crosscaps accordingly and keeping the cycles $C_{17k+1},\dots,C_{18k},$ we obtain a mixed surface grid of order $k$ as desired. See Figure 7 for an illustration.
@@
Let $H$ be a mixed surface grid of order $162^{h_{0}+c+h}\cdot k$ with $h$ handles and $c$ crosscaps.
For each of the $c$ crosscaps we need to call Subsection 2.2 at most $h$ times to achieve a situation where all handles are consecutive and thus obtain a surface grid. That is, if we aim for a surface grid of order $w$ with $h$ handles and $c$ crosscaps, we may start out with a mixed surface grid of order $9^{c+h}\cdot w.$ Moreover, this is enough to move any crosscap essentially anywhere along the cycle $C_{1}.$ That means that also the surface grid of order $9^{c+h}\cdot w$ contains every mixed surface grid of order $w$ with $h$ handles and $c$ crosscaps as a minor.
In particular, if $H^{\prime}$ is the surface grid of order $18^{h_{0}}\cdot k,$ then $H^{\prime}$ is a minor of $H$ by the reasoning above.
Now, to obtain a Dyck-grid of order $w$ with $h+h_{0}$ handles and $c_{0}$ crosscaps from a surface grid with $h$ handles and $c$ crosscaps we may proceed as follows: Choose three consecutive crosscaps at positions $i,$ $i+1,$ and $i+2$ such that the transaction at position $i-1$ is either a handle, or $i-1=0.$ Then apply Subsection 2.3 to these three crosscaps. The resulting minor we find is still a surface grid, but now on $h+1$ handles and $c-2$ crosscaps and its order has decreased by a factor of $18$ As we have to apply this exactly $h_{0}$ times to reduce the number of crosscaps below three, by starting out with a surface grid of order $18^{h_{0}}\cdot w$ with $h$ handles and $c$ crosscaps, we obtain the Dyck-grid of order $w$ and Euler-genus $2(h+h_{0})+c_{0}=2h+c$ as a minor. Notice that Subsection 2.3 states that reversing this operation also costs a factor of $18,$ that means that the (orientable in case $c=0$ and otherwise non-orientable) Dyck-grid of order $18^{h_{0}}\cdot w$ contains the surface grid of order $w$ with $h$ handles and $c$ crosscaps as a minor. In fact, for any $h^{\prime}\leq h_{0},$ it also contains the surface grid of order $w$ with $h+(h_{0}-h^{\prime})$ handles and $c-2(h_{0}-h^{\prime})$ crosscaps as a minor.
In particular, our graph $H^{\prime}$ from above contains a Dyck-grid of order $k$ and Euler-genus $2h+c$ as a minor.
Now let $g\coloneqq 2h+c$ as in the statement of our theorem and notice the following inequalities:
$c$ $\leq g,$
$2h$ $\leq g,\text{ and}$
$2h_{0}$ $\leq c,$
@@
$162^{h_{0}+c+h}\cdot k$ $\leq 162^{\frac{1}{2}g+g+\frac{1}{2}g}\cdot k$
@@
This is true for any choice of $h^{\prime}$ and $c^{\prime}$ such that $2h^{\prime}+c^{\prime}=g.$ Hence, by the discussion above, that any mixed surface grid of order $162^{2g}\cdot k$ with $h^{\prime}$ handles and $c^{\prime}$ crosscaps where $2h^{\prime}+c^{\prime}=g$ contains a Dyck-grid of order $k$ and Euler-genus $g$ as a minor. Moreover, the (orientable in case $c^{\prime}=0$ and otherwise non-orientable) Dyck-grid of order $162^{2g}k$ contains every mixed surface grid of order $k$ with $h^{\prime}$ handles and $c^{\prime}$ crosscaps as a minor. Hence, our proof is complete. ∎
@@ 3 A brief introduction to the toolset of graph minors
To prove the upper bound in Theorem 1.2, that is ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}\preceq{\mathsf{sobs}(\mathbb{S})}\mbox{-}{\sf bg},$ we make use of the powerful tools developed around the Graph Minors Series of Robertson and Seymour. In particular, we will be using a theorem developed by Kawarabayashi, Thomas, and Wollan in their proof of the GMST [32] which makes the extraction of some surface grid a very achievable task. However, to be able to even state this result we need to introduce a lot of machinery. The purpose of this section is the introduction and collection of the various technical definitions from [32] we need.
3.1 Walls an tangles
@@
An $(n\times m)$-grid is the graph $G_{n,m}$ with vertex set $[n]\times[m]$ and edge set
@@
We call the path $(i,1)(i,2)\dots(i,m)$ the $i$th row and the path $(1,j)(2,j)\dots(n,j)$ the $j$th column of the grid. An elementary $k$-wall $W_{k}$ for ${k\geq 3},$ is obtained from the ${(k\times 2k)}$-grid $G_{k,2k}$ by deleting every odd edge in every odd column and every even edge in every even column, and then deleting all degree-one vertices. The rows of $W_{k}$ are the subgraphs of $W_{k}$ induced by the rows of $G_{k,2k},$ while the $j$th column of $W_{k}$ is the subgraph induced by the vertices of columns ${2j-1}$ and ${2j}$ of $G_{k,2k}.$ We define the perimeter of $W_{k}$ to be the subgraph induced by ${\{(i,j)\in V(W_{k})\mid j\in\{1,2,2k,2k-1\}\text{ and }i\in[k]\textnormal{, or }i\in\{1,k\}\text{ and }j\in[2k]\}}.$ A $k$-wall $W$ is a graph isomorphic to a subdivision of $W_{k}.$ The vertices of degree three in $W$ are called the branch vertices. In other words, $W$ is obtained from a graph $W^{\prime}$ isomorphic to $W_{k}$ by subdividing each edge of $W^{\prime}$ an arbitrary (possibly zero) number of times. The perimeter of $W,$ denoted by ${\sf Perimeter}{W},$ is the subgraph isomorphic to the subgraph of $W^{\prime}$ induced by the vertices of the perimeter of $W_{k}$ together with the subdivision vertices of the edges of the perimeter of $W_{k}.$ We define rows and columns of $k$-walls analogously to their definition for elementary walls. A wall is a $k$-wall for some $k.$
@@
There exists a function $\mathsf{p}\colon\mathbb{N}\to\mathbb{N}$ with $\mathsf{p}(n)\in\mathcal{O}(n^{9}\operatorname{polylog}(n))$ such that for every $k\in\mathbb{N}$ and every graph $G,$ if ${\sf tw}(G)>\mathsf{p}(k)$ then $G$ contains a $k$-wall as a subgraph.
@@
Similarly, let $h\in\mathbb{N}$ and $c\in[0,2]$ and let $D$ be an $(h,c;t)$-Dyck-wall in $G$. Then also $D$ defines a majority side for any separation of order less than $r$. In this case the majority side is the sida that contains a full cycle and a column of the annulus wall of $D$. In the same way as above this allows us to define the tangle $\mathcal{T}_{D}$ induced by $D$.
@@
Let $α\in[2/3,1),$ $G$ be a graph and $S\subseteq V(G)$ be a vertex set. A set $X\subseteq V(G)$ is said to be an $α$-balanced separator for $S$ if for every component $C$ of $G-X$ it holds that $|C(C)\cap S|\leq α|S|.$ Let $q\in\mathbb{N}.$ We say that $S$ is a $(q,α)$-well-linked set of $G$ if there is no $α$-balanced separator of size at most $q$ for $S$ in $G.$ Given a $(q,α)$ well-linked set of $G$ we define
$\mathcal{T}_{S}\coloneqq\{(A,B)\in\mathcal{S}^{q+1}\mid|S\cap B|>α|S|\}.$
@@
Since it is easier to represent well-linked sets of obunded size with a bounded amount of memory, we opt to work with well-linked sets instead of tangles throughout this paper. Please note that, with the observation above, this still allows us to make use of all results of graph minors that need a large oder tangle as part of the input.
@@
Let $G$ and $H$ be graphs and let $p\coloneqq|V(H)|.$ Let $\mathcal{T}$ be a tangle in $G.$ An $H$-expansion in $G$ is controlled by $\mathcal{T}$ if for every separation $(A,B)\in\mathcal{T}$ with $|A\cap B|<p,$ there does not exist a branch vertex of the expansion contained in $A\setminus B.$ For a set $Z\subseteq V(G)$ we denote by $\mathcal{T}-Z$ the tangle in $G-Z$ formed by the separations $(A,B)$ in $G-Z$ such that $(A\cup Z,B\cup Z)\in\mathcal{T}.$ Given a surface $\Sigma$ and a $\Sigma$-decomposition $\delta$ of $G,$ we say that the decomposition is $\mathcal{T}$-central if for every separation $(A,B)\in\mathcal{T},$ there does not exist a cell $c\in C(\delta)$ such that $B\subseteq V(\Sigma_{\delta}(c)).$
Next we need a tool that allows us to relate a given $\Sigma$-decomposition to some tangle. We add a small bit of information to the statement of the lemma which is part of the proof of Lemma 15.3 in [32] but not made explicit in their formulation of the statement.
A tangle $\mathcal{T}^{\prime}$ in $G$ of order at most $k$ is a restriction of $\mathcal{T}$ if $\mathcal{T}^{\prime}\subseteq\mathcal{T}$.
[…]
Let $G$ be a graph, and let $k,\ell,r\geq 0$ such that $r\geq k(2\ell+1)+6$. Let $W$ be an $r$-wall in $G$. Let $\mathcal{T}$ be a tangle of order at least $k(2\ell+1)+6$ such that $\mathcal{T}_{W}$ is a restriction of $\mathcal{T}$. Assume that there exists a $\Sigma$-decomposition $\delta$ of $G$ of breadth $k$ and depth $\ell$ such that $W$ is flat in $\delta$. Then there exists $Z\subseteq V(G),$ $|Z|\leq k(2\ell+1)$ and a $\Sigma$-decomposition $\delta^{\prime}$ of $G-Z$ of breadth at most $k$ and depth at most $\ell$ which is $(\mathcal{T}-Z)$-central.
Moreover, the set $Z$ consists of at most $12\ell+$ 1 vertices from each vortex of $\delta$ and every vertex of $Z$ belongs of some vortex of $\delta^{\prime}$.

### v4

@@ 3.1 A graph-minor analogue of Dyck’s lemma
Let $k,\mathsf{c},\mathsf{h}\geq 0$ be integers such that $k\geq 1$ and $c\geq 3.$ Let $H^{\prime}$ be the $(18k,4(\mathsf{c}+\mathsf{h}+1)\cdot 18k)$-cylindrical grid and $H^{\prime\prime}$ be the $(18k,4(c-2+\mathsf{h}+1+1)\cdot 18k)$-cylindrical grid. Let $I^{\prime}_{\mathsf{c}},I^{\prime}_{\mathsf{h}}$ be a partition of $[2,\mathsf{c}+\mathsf{h}+1]$ where $|I^{\prime}_{\mathsf{c}}|=c$ and $|I^{\prime}_{\mathsf{h}}|=h$ and there exists some integer $i\in I^{\prime}_{\mathsf{c}}$ such that $i+1,i+2\in I^{\prime}_{\mathsf{c}},$ moreover, let $I^{\prime\prime}_{\mathsf{c}},I^{\prime\prime}_{\mathsf{h}}$ be a partition of $[2,\mathsf{c}+\mathsf{h}]$ such that $i\in I^{\prime\prime}_{\mathsf{h}}$ and $i+1\in I^{\prime\prime}_{\mathsf{c}}.$ Finally, let $H^{1}$ be the $(k,4(\mathsf{c}+\mathsf{h})\cdot k)$-cylindrical grid and $H^{2}$ be the $(k,4(\mathsf{c}+\mathsf{h}+1)\cdot k)$-cylindrical grid.
@@
    the mixed surface grid $H^{\prime}_{I^{\prime}_{\mathsf{c}},I^{\prime}_{\mathsf{h}}}$ of order $18k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles contains the mixed surface grid $H^{1}_{I^{\prime\prime}_{\mathsf{c}},I^{\prime\prime}_{\mathsf{h}}}$ of order $k$ with $\mathsf{c}-2$ crosscaps and $\mathsf{h}+1$ handles as a minor, and
@@
    the mixed surface grid $H^{\prime\prime}_{I^{\prime\prime}_{\mathsf{c}},I^{\prime\prime}_{\mathsf{h}}}$ of order $18k$ with $\mathsf{c}-2$ crosscaps and $\mathsf{h}+1$ handles contains the mixed surface grid $H^{2}_{I^{\prime}_{\mathsf{c}},I^{\prime}_{\mathsf{h}}}$ of order $k$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles as a minor.
@@
The final step is to untangle the crosscap at position $i+2$ from the newly created handle. Notice that “position $i+2$” refers only to the bundle of paths forming the crosscap transactions as we are not in a mixed surface grid any more. Indeed, by merging the other two crosscaps into a single handle, we have lost a position. Similar to the proof of Section 3 we fix two disjoint intervals of endpoints, one within the right most part of the crosscap transaction on the cycle $C_{18k},$ and the other in the right most part of the middle transaction that is part of the handle. Both intervals are of order $2k.$ We then route a crosscap transaction from the left of these intervals to the right. As depicted in Figure 5 we may traverse the handle transaction which allows us to traverse the crosscap exactly once. Hence, the resulting transaction is, indeed, a crosscap. Since each of the two transactions involved in the handle are of order $12k,$ and we only spend $4k$ of their paths to route the crosscap, there are enough paths left to choose $k$ paths each in order to form a handle transaction of order $k.$ By restricting all other handles and crosscaps accordingly and keeping the cycles $C_{17k+1},\dots,C_{18k},$ we obtain a mixed surface grid of order $k$ as desired. See Figure 6 for an illustration.
@@
Let $H$ be a mixed surface grid of order $162^{\mathsf{h}_{0}+\mathsf{c}+\mathsf{h}}\cdot k$ with $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps.
For each of the $\mathsf{c}$ crosscaps we need to call Section 3 at most $\mathsf{h}$ times to achieve a situation where all handles are consecutive and thus obtain a surface grid. That is, if we aim for a surface grid of order $w$ with $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps, we may start out with a mixed surface grid of order $9^{\mathsf{c}+\mathsf{h}}\cdot w.$ Moreover, this is enough to move any crosscap essentially anywhere along the cycle $C_{1}.$ That means that also the surface grid of order $9^{\mathsf{c}+\mathsf{h}}\cdot w$ contains every mixed surface grid of order $w$ with $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps as a minor.
In particular, if $H^{\prime}$ is the surface grid of order $18^{\mathsf{h}_{0}}\cdot k,$ then $H^{\prime}$ is a minor of $H$ by the reasoning above.
Now, to obtain a Dyck-grid of order $w$ with $\mathsf{h}+\mathsf{h}_{0}$ handles and $\mathsf{c}_{0}$ crosscaps from a surface grid with $h$ handles and $\mathsf{c}$ crosscaps we may proceed as follows: Choose three consecutive crosscaps at positions $i,$ $i+1,$ and $i+2$ such that the transaction at position $i-1$ is either a handle, or $i-1=0.$ Then apply Subsection 3.1 to these three crosscaps. The resulting minor we find is still a surface grid, but now on $\mathsf{h}+1$ handles and $\mathsf{c}-2$ crosscaps and its order has decreased by a factor of $18$ As we have to apply this exactly $\mathsf{h}_{0}$ times to reduce the number of crosscaps below three, by starting out with a surface grid of order $18^{\mathsf{h}_{0}}\cdot w$ with $\mathsf{h}$ handles and $\mathsf{c}$ crosscaps, we obtain the Dyck-grid of order $w$ and Euler-genus $2(\mathsf{h}+\mathsf{h}_{0})+\mathsf{c}_{0}=2\mathsf{h}+\mathsf{c}$ as a minor. Notice that Subsection 3.1 states that reversing this operation also costs a factor of $18,$ that means that the (orientable in case $\mathsf{c}=0$ and otherwise non-orientable) Dyck-grid of order $18^{\mathsf{h}_{0}}\cdot w$ contains the surface grid of order $w$ with $h$ handles and $\mathsf{c}$ crosscaps as a minor. In fact, for any $\mathsf{h}^{\prime}\leq\mathsf{h}_{0},$ it also contains the surface grid of order $w$ with $\mathsf{h}+(\mathsf{h}_{0}-\mathsf{h}^{\prime})$ handles and $\mathsf{c}-2(\mathsf{h}_{0}-\mathsf{h}^{\prime})$ crosscaps as a minor.
In particular, our graph $H^{\prime}$ from above contains a Dyck-grid of order $k$ and Euler-genus $2\mathsf{h}+\mathsf{c}$ as a minor.
Now let $g\coloneqq 2\mathsf{h}+\mathsf{c}$ as in the statement of our theorem and notice the following inequalities:
$\mathsf{c}$ $\leq g,$
$2\mathsf{h}$ $\leq g,\text{ and}$
$2\mathsf{h}_{0}$ $\leq\mathsf{c},$
@@
$162^{\mathsf{h}_{0}+\mathsf{c}+\mathsf{h}}\cdot k$ $\leq 162^{\frac{1}{2}g+g+\frac{1}{2}g}\cdot k$
@@
This is true for any choice of $\mathsf{h}^{\prime}$ and $\mathsf{c}^{\prime}$ such that $2\mathsf{h}^{\prime}+\mathsf{c}^{\prime}=g.$ Hence, by the discussion above, that any mixed surface grid of order $162^{2g}\cdot k$ with $\mathsf{h}^{\prime}$ handles and $\mathsf{c}^{\prime}$ crosscaps where $2\mathsf{h}^{\prime}+\mathsf{c}^{\prime}=g$ contains a Dyck-grid of order $k$ and Euler-genus $g$ as a minor. Moreover, the (orientable in case $\mathsf{c}^{\prime}=0$ and otherwise non-orientable) Dyck-grid of order $162^{2g}k$ contains every mixed surface grid of order $k$ with $\mathsf{h}^{\prime}$ handles and $\mathsf{c}^{\prime}$ crosscaps as a minor. Hence, our proof is complete. ∎
3.3 Universal graphs for graphs of bounded Euler-genus.
An important consequence of Theorem 3.1 is that Dyck-walls are indeed minor-universal for graphs of bounded Euler-genus. Indeed, any mixed surface wall with $h$ handles and $\mathsf{c}$ crosscaps is minor-universal for graphs that embed into the surface obtained from the sphere by adding $h$ handles and $\mathsf{c}$ crosscaps.
Let us now state the theorem of Gavoille and Hilaire.
[…]
For every pair $(\mathsf{h},\mathsf{c})\in\mathbb{N}\times[0,2],$ a graph $G$ is embeddable in $\Sigma^{(\mathsf{h},\mathsf{c})}$ if and only if it is a there exists a $k=k(G)$ such that one of the following holds
  • in case $\mathsf{c}=0,$ $G$ is a minor of $\mathscr{D}^{(\mathsf{h},0)}_{k},$ and
  • in case $\mathsf{c}\neq 0,$ $G$ is a minor of the mixed surface grid of order $k$ with 00 handles and $2\mathsf{h}+\mathsf{c}$ crosscaps.
Moreover, we have that $k(G)\in\mathcal{O}((2\mathsf{h}+\mathsf{c})^{2}(2\mathsf{h}+\mathsf{c}+|V(G)|)^{2}).$
We now obtain the fact that Dyck-walls are minor-universal for graphs of bounded Euler-genus as a direct corollary of Subsection 3.3 and Theorem 3.1.
[…]
For every pair $(\mathsf{h},\mathsf{c})\in\mathbb{N}\times[0,2],$ a graph $G$ is embeddable in $\Sigma^{(\mathsf{h},\mathsf{c})}$ if and only if it is a minor of $\mathscr{D}^{(\mathsf{h},\mathsf{c})}_{k},$ for some $k=k(G)\in\mathbb{N}.$ Moreover, $k(G)\in 2^{\mathcal{O}(\mathsf{h}+\mathsf{c})}.$
@@ 4 A brief introduction to the toolset of graph minors
We will be using a theorem developed by Kawarabayashi, Thomas, and Wollan in their proof of the GMST [17] which makes the extraction of some surface grid a very achievable task. However, to be able to even state this result we need to introduce a lot of machinery. The purpose of this section is the introduction and collection of the various technical definitions from [17] we need.
4.1 Walls and tangles
@@
An $(n\times m)$-grid is the graph $\Gamma_{n,m}$ with vertex set $[n]\times[m]$ and edge set
@@
We call the path $(i,1)(i,2)\dots(i,m)$ the $i$th row and the path $(1,j)(2,j)\dots(n,j)$ the $j$th column of the grid. An elementary $k$-wall $W_{k}$ for ${k\geq 3},$ is obtained from the ${(k\times 2k)}$-grid $\Gamma_{k,2k}$ by deleting every odd edge in every odd column and every even edge in every even column, and then deleting all degree-one vertices. The rows of $W_{k}$ are the subgraphs of $W_{k}$ induced by the rows of $\Gamma_{k,2k},$ while the $j$th column of $W_{k}$ is the subgraph induced by the vertices of columns ${2j-1}$ and ${2j}$ of $\Gamma_{k,2k}.$ We define the perimeter of $W_{k}$ to be the subgraph induced by ${\{(i,j)\in V(W_{k})\mid j\in\{1,2,2k,2k-1\}\text{ and }i\in[k]\textnormal{, or }i\in\{1,k\}\text{ and }j\in[2k]\}}.$ A $k$-wall $W$ is a graph isomorphic to a subdivision of $W_{k}.$ The vertices of degree three in $W$ are called the branch vertices. In other words, $W$ is obtained from a graph $W^{\prime}$ isomorphic to $W_{k}$ by subdividing each edge of $W^{\prime}$ an arbitrary (possibly zero) number of times. The perimeter of $W,$ denoted by ${\sf Perimeter}(W),$ is the subgraph isomorphic to the subgraph of $W^{\prime}$ induced by the vertices of the perimeter of $W_{k}$ together with the subdivision vertices of the edges of the perimeter of $W_{k}.$ We define rows and columns of $k$-walls analogously to their definition for elementary walls. A wall is a $k$-wall for some $k.$
@@
There exists a function $\mathsf{p}\colon\mathbb{N}\to\mathbb{N}$ with $\mathsf{p}(n)\in\mathcal{O}(n^{9}\operatorname{\mathsf{polylog}}(n))$ such that for every $k\in\mathbb{N}$ and every graph $G,$ if ${\sf tw}(G)>\mathsf{p}(k)$ then $G$ contains a $k$-wall as a subgraph.
@@
Let $\mathcal{T}$ be a tangle in a graph $G.$ A tangle $\mathcal{T}^{\prime}$ in $G$ is a truncation of $\mathcal{T}$ if $\mathcal{T}^{\prime}\subseteq\mathcal{T}.$
@@
Similarly, let $h\in\mathbb{N}$ and $c\in[0,2]$ and let $D$ be an $(\mathsf{c},\mathsf{h};t)$-Dyck-wall in $G.$ Then also $D$ defines a majority side for any separation of order less than $r.$ In this case the majority side is the side that contains a full cycle and a column of the annulus wall of $D.$ In the same way as above this allows us to define the tangle $\mathcal{T}_{D}$ induced by $D.$
@@
Let $\alpha\in[2/3,1),$ $G$ be a graph and $S\subseteq V(G)$ be a vertex set. A set $X\subseteq V(G)$ is said to be an $\alpha$-balanced separator for $S$ if for every component $C$ of $G-X$ it holds that $|C(C)\cap S|\leq\alpha|S|.$ Let $q\in\mathbb{N}.$ We say that $S$ is a $(q,\alpha)$-well-linked set of $G$ if there is no $\alpha$-balanced separator of size at most $q$ for $S$ in $G.$ Given a $(q,\alpha)$ well-linked set of $G$ we define
$\mathcal{T}_{S}\coloneqq\{(A,B)\in\mathcal{S}^{q+1}\mid|S\cap B|>\alpha|S|\}.$
@@
Since it is easier to represent well-linked sets of unbounded size with a bounded amount of memory, we opt to work with well-linked sets instead of tangles throughout this paper. Please note that, with the observation above, this still allows us to make use of all results of graph minors that need a large order tangle as part of the input.